a 4kg block of wood is pulled along a horizontal ground from rest and a force of 15 N is required to produce an acceleration of 2 m/s^2. In order to pull the block of wood with constant speed 5 m/s on the same horizontal ground. what should of the magnitude of the force be?
1
Expert's answer
2018-11-19T16:03:10-0500

F^'=F-ma

F^'=15-(4)(2)=7 N.